Show complete solution

1.f(x)=2 sin (x-​π/​4)

2.f(x)=sin (x+π/​3)

3.f(x)=3 cos (x+π/​6)

4.f(x)=2 cos (x-2π)

5.f(x)=tan x


1
Expert's answer
2021-03-24T07:44:00-0400

1.f(x)=2 sin (x-​π/​4)

Solution: x-​π/​4=​πk, k€Z
x=​π/​4+​πk, k€Z

2.f(x)=sin (x+π/​3)

Solution :x+π/​3=​πk, k€Z
x=-π/​3+​πk, k€Z

3.f(x)=3 cos (x+π/​6)

Solution: x+π/​6=π/​2+​πk, k€Z
x=-π/​6+π/​2+​πk, k€Z
x=π/​3+​πk, k€Z

4.f(x)=2 cos (x-2π)

Solution:cos(x-2π)=cos(x) 
x=π/​2+​πk, k€Z

5.f(x)=tan x

Solution: x=πk, k€Z